# The Cosmological CPT Theorem

### v2

@@ 8 Discussion
There are a number of other directions in which this work can be usefully extended. Our results lead fairly directly to a no-go theorem for certain specific cosmological correlators of bosonic fields [88] derived from the phase formula in this paper. An interesting direction which should certainly be explored is using a similar formalism to constrain spinors using discrete symmetries, which is yet to be done in the cosmology literature. To derive such constraints it will be necessary for us to use projective representations of the group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ with the appropriate double cover of the group for half-integer spin fields being the dihedral group $D_{4}$. Further work should also certainly be done on extending the analysis of the boundary wavefunction coefficients to the case of heavy fields in the principal series (discussed by [76, 77, 78, 79, 80, 81, 82, 83, 84, 85]). In this case, the two dimensions $\Delta=d/2\pm i\nu$ correspond to equally-leading and oscillatory modes as $\eta\to 0^{-}$, leaving to subtle questions about the best way to define the boundary conditions at ${\cal I}^{+}$. Presumably, the same analytic continuation in $\eta$ which led to the phase rule, will now lead to an enhancement of one mode over the other by an amount exponential in $\nu$.
@@ Acknowledgements
It is a pleasure to thank Santiago Agüí Salcedo, Gonçalo Araújo Regado, Dionysios Anninos, Tarek Anous, Nima Arkani-Hamed, Paolo Benincasa, Alejandra Castro, Calvin Chen, Chandramouli Chowdhury, Nick Cooper, Mang Hei Gordon Lee, Hofie Sigridar Hannesdottir, Daniel Harlow, Ted Jacobson, Rohit Kalloor, Rifath Khan, Shota Komatsu, Juan Maldacena, Ciaran McCulloch, Prahar Mitra, Sebastian Mizera, Enrico Pajer, Guilherme Pimentel, Bharathkumar Radhakrishnan, Adel Rahman, Muthusamy Rajaguru, Alan Rios Fukelman, Vladimir Schaub, Vasudev Shyam, David Stefanyszyn, Leonard Susskind, Marija Tomašević, Xi Tong, Erik Verlinde, Tom Westerdijk, Edward Witten, and Yuhang Zhu for useful discussions. AT would like to especially emphasise the importance of the initial collaboration with Carlos Duaso Pueyo, who was the first to encourage him to continue pursuing this research program.
HG was supported by the Science and Technology Facilities Council through a postgraduate studentship and the Cambridge Trust Vice Chancellor’s Award, and is supported by a Postdoctoral Fellowship at National Taiwan University funded by the Ministry of Education (MOE) NTU-112L4000-1. AT was supported in part by the Heising-Simons Foundation, the Simons Foundation, the Bell Burnell Graduate Scholarship Fund, the Cavendish (University of Cambridge) and a KITP Graduate fellowship. AT and AW are supported by the AFOSR grant FA9550-19-1-0260 “Tensor Networks and Holographic Spacetime” and grants no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). AW was also supported by the STFC grant ST/P000681/1 “Particles, Fields and Extended Objects”, an Isaac Newton Trust Early Career grant. AW was also partly supported by NSF grant PHY-2207584 while finishing this paper during his sabbatical at the IAS.

### v3

@@ 8 Discussion
Our phase rule has important implications for the dS/CFT correspondence, providing a new tool to determine the phases of correlators in the holographic dual conformal field theory, thereby advancing our understanding of quantum gravity in cosmological settings. As dS/CFT theories are non-unitary—in the sense that the boundary theory, thought of as a Euclidean field theory, does not itself satisfy a boundary-$\bf{RR}$, they do not satisfy the reality conditions of normal unitary theories. It is therefore critical to identify the (presumably equally constraining) properties that a dS/CFT must satisfy to be dual to a unitary bulk. This article provides substantial progress towards this goal, but further work is needed to clarify what this constraint means from the perspective of the boundary field theory. It would also be useful if this reality constraint can be promoted to a more restrictive positivity constraint on dS/CFT, similar to how Reflection Positivity is more restrictive than Reflection Reality in unitary theories. Put another way, our reality conditions are necessary for bulk unitarity, but it would be even nicer to identify a necessary and sufficient condition for bulk unitarity.
There are a number of other directions in which this work can be usefully extended. Our results lead fairly directly to a no-go theorem for certain specific cosmological correlators of bosonic fields [88] derived from the phase formula in this paper. In general, a constraint on the phase of a wavefunction coefficient is not usually easy to observe, because it requires looking at correlators involving time derivatives $\pi$, which are subleading compared to correlators involving the field $\phi$ itself. It is only at special values of $\Delta$, e.g. the cases addressed by [88], that there is a no-go result on certain correlators becoming large. For generic values of $\Delta$, it is more feasible to view our phase rule as a way to determine the difficult-to-observe (imaginary) part of $\log\Psi$, based on the more easily observable (real) part of $\log\Psi$.
@@ Acknowledgements
It is a pleasure to thank Santiago Agüí Salcedo, Gonçalo Araújo Regado, Dionysios Anninos, Tarek Anous, Nima Arkani-Hamed, Paolo Benincasa, Alejandra Castro, Calvin Chen, Chandramouli Chowdhury, Nick Cooper, Mang Hei Gordon Lee, Hofie Sigridar Hannesdottir, Daniel Harlow, Ted Jacobson, Rohit Kalloor, Rifath Khan, Shota Komatsu, Juan Maldacena, Ciaran McCulloch, Prahar Mitra, Sebastian Mizera, Enrico Pajer, Guilherme Pimentel, Bharathkumar Radhakrishnan, Adel Rahman, Muthusamy Rajaguru, Alan Rios Fukelman, Vladimir Schaub, Vasudev Shyam, David Stefanyszyn, Leonard Susskind, Marija Tomašević, Xi Tong, Erik Verlinde, Tom Westerdijk, Edward Witten, and Yuhang Zhu for useful discussions. AT would like to especially emphasise the importance of the initial collaboration with Carlos Duaso Pueyo, who was the first to encourage him to continue pursuing this research program. We are also grateful to an anonymous reviewer, who provided helpful comments on an earlier draft of the paper.
HG was supported by the Science and Technology Facilities Council through a postgraduate studentship and the Cambridge Trust Vice Chancellor’s Award, and is supported by a Postdoctoral Fellowship at National Taiwan University funded by the Ministry of Education (MOE) NTU-112L4000-1. AT was supported in part by the Heising-Simons Foundation, the Simons Foundation, the Bell Burnell Graduate Scholarship Fund, the Cavendish (University of Cambridge) and a KITP Graduate fellowship. AT and AW are supported by the AFOSR grant FA9550-19-1-0260 “Tensor Networks and Holographic Spacetime” and grants no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). AW was also supported by the STFC grants ST/P000681/1 “Particles, Fields and Extended Objects”, ST/X000664/1 “Quantum Fields, Quantum Gravity and Quantum Particles”, an Isaac Newton Trust Early Career grant. AW was also partly supported by NSF grant PHY-2207584 while finishing this paper during his sabbatical at the IAS.